# MOCHa: Multi-Objective Reinforcement Mitigating Caption Hallucinations

### v1

@@ Abstract
While recent years have seen rapid progress in image-conditioned text generation, image captioning still suffers from the fundamental issue of hallucinations, the generation of spurious details that cannot be inferred from the given image. Dedicated methods for reducing hallucinations in image captioning largely focus on closed-vocabulary object tokens, ignoring most types of hallucinations that occur in practice. In this work, we propose MOCHa, an approach that harnesses advancements in reinforcement learning (RL) to address the sequence-level nature of hallucinations in an open-world setup. To optimize for caption fidelity to the input image, we leverage ground-truth reference captions as proxies to measure the logical consistency of generated captions. However, optimizing for caption fidelity alone fails to preserve the semantic adequacy of generations; therefore, we propose a multi-objective reward function that jointly targets these qualities, without requiring any strong supervision. We demonstrate that these goals can be simultaneously optimized with our framework, enhancing performance for various captioning models of different scales. Our qualitative and quantitative results demonstrate MOCHa’s superior performance across various established metrics. We also demonstrate the benefit of our method in the open-vocabulary setting. To this end, we contribute OpenCHAIR, a new benchmark for quantifying open-vocabulary hallucinations in image captioning models, constructed using generative foundation models. We will release our code, benchmark, and trained models.
BLIP +MOCHa
A pregnant mother fills coffee into her new
white cup
Coffee being poured into a white cup on a wooden table
Dimly shining coffee drink on top of wooden table with brown donut
A glass mug of coffee on a wooden table
Figure 1: Hallucinated details (shown as highlighted text) are prevalent in the outputs of modern image captioning models, as exemplified by generations sampled from BLIP [20] above. Our reinforcement learning-based approach adjusts captioning models to output detailed, valid captions while avoiding such hallucinations, as illustrated in the right-hand column (+MOCHa).
@@ 1 Introduction
Image captioning, the task of generating text that describes an image, is one of the most fundamental machine learning tasks combining vision and language. Unfortunately, hallucinations plague the current state-of-the-art (SOTA), making it less usable for practical tasks that require confidence in the factual correctness of generated captions. Consider, for instance, the images in Figure 1. SOTA image captioning models can generate text that is highly semantically related to its associated imagery, but also contains spurious details (“dimly shining”, “brown donut”). Such hallucinated spurious details either damage user confidence or lead to uncritical acceptance of fallacious (and even potentially dangerous) generated content [5, 25, 4].
@@
Conceptually, hallucinations in image captioning (and text generation in general) models stem from deficiencies in the standard language modeling (LM) objective. The token-level likelihood maximization LM objective does not directly optimize the sequence-level quality of generated text, and factual groundedness is inherently a sequence-level property of text. However, prior works mitigating hallucinations in image captioning mostly avoid the global sequence-level nature of hallucination by limiting their scope to a fixed set of possible object tokens (e.g. objects in MS-COCO) [2, 24, 28].
Figure 2: The MOCHa framework. The algorithm iteratively collects a batch of data from an image captioning model $M$ (left side) and then applies an optimization step to this model (right side). The multi-objective reward reinforces $M$ to produce captions closer to the high-scoring captions and further from the low-scoring captions.
In general, the inability of LM to optimize sequence-level properties has inspired the use of reinforcement learning (RL) as an additional fine-tuning step to align text generation models to desired behaviors. This has been applied to image captioning to optimize textual similarity to ground truth semantics [30, 35], and more recently, to align large language models (LLMs) and large vision-and-language models (VLMs) to human preferences [49, 36, 26]. However, existing applications of RL to NLP and multimodal learning focus on optimizing general text quality or estimated human preferences, which do not directly address fidelity to an input condition (e.g. consistency with an input image, in the case of image captioning) and hence, may lead to increased hallucinations in generated text [26].
@@
In this work, we introduce MOCHa (Multi-Objective reinforcement mitigating Caption Hallucinations) – a RL-based approach for mitigating image captioning hallucinations in an open-world setup. We observe that RL applied to caption fidelity alone fails to preserve the semantic adequacy (i.e. descriptiveness) of output text, while optimizing for the latter does not enforce factually grounded text. Our key insight is that these two goals can be jointly optimized at the sequence-level by applying RL with a multi-objective reward function. Furthermore, we perform this optimization fully automatically by leveraging SOTA text-based learned metrics, without requiring direct supervision.
Finally, while established benchmarks and metrics for quantifying hallucinations in captioning models exist for closed-vocabulary object sets, such benchmarks (to the best of our knowledge) do not exist in an open-vocabulary setup. Therefore, we introduce OpenCHAIR, a new benchmark for quantifying hallucinations over an open-vocabulary set of objects, constructed automatically using LLMs and powerful text-to-image generation models. We show that our approach can be flexibly applied to a variety of captioning architectures and sizes, and it improves performance over a wide set of existing metrics used by prior works and also over our new benchmark.
Explicitly stated, our key contributions are:
@@
  • The MOCHa framework for optimizing a wide array of VLMs to produce high-quality factually-grounded output.
  • The OpenCHAIR benchmark for evaluating open-vocabulary hallucinations in image captioning models.
  • Results showing that MOCHa outperforms existing methods for hallucination mitigation on open- and closed-vocabulary measures while preserving caption quality.

### v3

# Mitigating Open-Vocabulary Caption Hallucinations

@@ Abstract
While recent years have seen rapid progress in image-conditioned text generation, image captioning still suffers from the fundamental issue of hallucinations, namely, the generation of spurious details that cannot be inferred from the given image. Existing methods largely use closed-vocabulary object lists to mitigate or evaluate hallucinations in image captioning, ignoring the long-tailed nature of hallucinations that occur in practice. To this end, we propose a framework for addressing hallucinations in image captioning in the open-vocabulary setting. Our framework includes a new benchmark, OpenCHAIR, that leverages generative foundation models to evaluate open-vocabulary object hallucinations for image captioning, surpassing the popular and similarly-sized CHAIR benchmark in both diversity and accuracy. Furthermore, to mitigate open-vocabulary hallucinations without using a closed object list, we propose MOCHa, an approach harnessing advancements in reinforcement learning. Our multi-objective reward function explicitly targets the trade-off between fidelity and adequacy in generations without requiring any strong supervision. MOCHa improves a large variety of image captioning models, as captured by our OpenCHAIR benchmark and other existing metrics. We will release our code and models.
Figure 1: Hallucinated details (shown as highlighted text) are prevalent in the outputs of modern image captioning models, such as the above generation sampled from BLIP2 Li et al. (2023a). By considering hallucinations in the open-vocabulary setting, we can both quantify and mitigate their effects, illustrated by the improvement provided by our RL-based MOCHa framework (+MOCHa).
Mitigating Open-Vocabulary Caption Hallucinations
Assaf Ben-Kish    Moran Yanuka    Morris Alper    Raja Giryes    Hadar Averbuch-Elor Tel-Aviv University https://assafbk.github.io/mocha
@@ 1 Introduction
Image captioning, the task of generating text that describes an image, is one of the most fundamental machine learning tasks combining vision and language. Unfortunately, hallucinations plague the current state-of-the-art (SOTA), making it less usable for practical tasks that require confidence in the factual correctness of generated captions. Consider, for instance, the image in Figure 1. SOTA image captioning models can generate text that is highly semantically related to its associated imagery, but also contains spurious details (“skateboard”). Such hallucinated spurious details either damage user confidence or lead to uncritical acceptance of fallacious (and even potentially dangerous) generated content Chong et al. (2022); McGowan et al. (2023); Chong et al. (2023).
Figure 2: The OpenCHAIR Benchmark. We illustrate the construction of the OpenCHAIR benchmark via an LLM and text-to-image generation model, and its usage for evaluating image captioning models. We first use captions from MS-COCO as seeds to generate diverse synthetic captions. Using syntactic parsing and filtering heuristics, we select for captions containing various open-vocabulary objects. We then generate images corresponding to these captions, producing our benchmark of images linked with object annotations. To evaluate a captioning model, we run it on this benchmark and compare predicted and GT object categories.
Hallucinations may take a variety of forms in text. However, prior work addressing hallucinations in image captioning has largely focused on detecting or mitigating hallucinations by using closed-vocabulary object lists. While this simplifies the problem under consideration, it fails to capture the diversity of hallucinations observed in modern image captioning models. Thus, we propose a framework for both quantifying and mitigating hallucinations in the open-vocabulary setting.
While established benchmarks and metrics for quantifying hallucinations in captioning models exist for closed-vocabulary object sets, they do not exist (to our knowledge) in an open-vocabulary setup. Accordingly, we introduce OpenCHAIR, a new benchmark for quantifying object hallucinations in an open-vocabulary setting. We construct our benchmark using text-to-image models and large language models (LLMs) for generating data and performing evaluation. This allows for capturing and accurately quantifying a wide variety of object hallucination types without being limited to a fixed set of categories. Moreover, our open-vocabulary evaluation method considers free-text predictions without referencing a fixed synonym list. Our evaluations show that this outperforms the CHAIR closed-vocabulary metric Rohrbach et al. (2018) at capturing performance over diverse hallucinations, providing a complementary measure to CHAIR’s evaluation over eighty common object types on natural images.
@@
Equipped with this metric, we turn to hallucination mitigation. A major cause for hallucinations in image captioning are deficiencies in the standard language modeling (LM) objective. The token-level language modeling objective does not directly optimize the sequence-level quality of generated text, and factual groundedness is inherently a sequence-level property of text. Yet, many prior works that directly optimize hallucinations in image captioning limit their scope to a fixed set of possible object tokens, e.g. objects in MS-COCO Biten et al. (2021); Liu et al. (2022); Petryk et al. (2023), which is incompatible with an open-vocabulary setting.
@@
To mitigate hallucinations without using a closed-vocabulary object list, we introduce MOCHa, a Multi-Objective reinforcement learning (RL) based approach for Mitigating Open-vocabulary Caption Hallucinations. We observe that RL applied to caption fidelity alone fails to preserve the semantic adequacy (i.e. descriptiveness) of output text, while optimizing for the latter does not enforce factually grounded text. Our key insight is that these two goals can be jointly optimized at the sequence-level by applying RL with a multi-objective reward function. Furthermore, we perform this optimization fully automatically by leveraging SOTA text-based learned metrics, without requiring direct supervision. By considering hallucinations in an open setting, we are able to improve performance across diverse hallucination types, as demonstrated by our OpenCHAIR benchmark as well as other metrics. Moreover, we show that our approach can be flexibly applied to a variety of captioning architectures and sizes.
@@
Explicitly stated, our key contributions are: (i) OpenCHAIR, a benchmark for open-vocabulary object hallucinations in image captioning. (ii) MOCHa, a framework for optimizing a wide array of VLMs to produce high-quality factually-grounded output. (iii) Experiments showing the advantage of OpenCHAIR for measuring hallucinations in the open setting, and of MOCHa for reducing them.
2 The OpenCHAIR Benchmark
[…]
In order to create a new benchmark that enables measuring the hallucination rate of arbitrary objects, while still maintaining high quality ground-truth captions, we use the pipeline illustrated in Figure 2. We first prompt the LLM Llama-2 Touvron et al. (2023) with few-shot examples of image captions from MS-COCO, having it generate captions with a similar style but containing diverse details (and in particular, objects that are likely not contained in the closed set of MS-COCO object labels). We then parse these synthetic captions with a syntactic parsing model, identify nouns with high concreteness scores Brysbaert et al. (2014) (as these generally represent concrete objects), and balance the generated captions among object types to cover a wide array of objects. Subsequently, we utilize the text-to-image diffusion model Stable Diffusion XL  Podell et al. (2023) to generate images from these newly formed captions. This process results in a dataset that consists of synthetic images with corresponding captions including diverse, open-vocabulary objects. While this approach naturally scales to any number of desired image-caption pairs, we generate 5K such pairs (the same order of items found in the widely-used MS-COCO Karpathy test split) and perform manual filtering to assure each pair’s alignment and general quality. In total, we removed a small minority (3%) of generated image-caption pairs. Figure 3 shows examples of image-captions pairs from OpenCHAIR.
Captioning models may predict free-text objects semantically matching the ground-truth while taking a different surface form (e.g. chihuahua vs. dog). To capture this in the open-vocabulary setting (rather than using a fixed list of synonyms as done in CHAIR), we evaluate captioning models as follows: After predicting a caption for each image in the OpenCHAIR dataset, we parse them to identify objects as described above. For each extracted object $o$, we compare it to the ground-truth synthetic caption $c$ by prompting an LLM, asking it whether an image with caption $c$ contains the object $o$ and using its answers to count hallucinations. Following CHAIR, we calculate the hallucination rate as $n_{h}/n_{tot}$, where $n_{h}$ is the number of hallucinated objects (no answers) and $n_{tot}$ is the total number of objects considered. Figure 4 illustrates the difference between OpenCHAIR evaluation and the closed-vocabulary CHAIR metric.
“A green emerald is perched on a rock in a cave."
“A group of mushrooms in the forest."
“A dog dressed as a human with a wig and eyeglasses."
Figure 3: OpenCHAIR Examples. We show examples of images from the OpenCHAIR benchmark along with their accompanying ground-truth captions, illustrating its diverse coverage of object types. Long captions are truncated due to space considerations.
3 The MOCHa Framework
To mitigate captioning hallucinations in the open-vocabulary setting, we propose MOCHa, an RL-based pipeline using SOTA methods for stable reinforcement along with a carefully designed reward function that jointly optimizes for caption fidelity and semantic adequacy. Figure 5 presents it. We turn to describe the learning procedure and objectives used in MOCHa. We start with preliminaries, then describe the reward function that MOCHa optimizes (Section 3.1), and finally present the RL algorithm used for optimization (Section 3.2).
Figure 4: OpenCHAIR vs. CHAIR. In the above the predicted object guitar would not be counted by CHAIR since it is not in its fixed vocabulary, while man would not be classified as a hallucination since it is defined by CHAIR as a synonym of child. In contrast, OpenCHAIR’s LLM classifies both as hallucinations.
Figure 5: MOCHa scheme. The algorithm iteratively collects a minibatch of data from an image captioning model $M$ (left side) and then applies an optimization step to the captioning model (right side). The multi-objective reward reinforces $M$ to produce captions closer to the high-scoring captions and further from the low-scoring captions.
Preliminaries. In general, RL views a model as an agent that interacts with the external environment and receives a reward, learning to optimize for this reward via exploring the environment Sutton and Barto (2018). In the case of image captioning, this model is a VLM operating in an environment of images and reference captions Rennie et al. (2017). During training, the agent generates a caption by sampling from its own predicted distribution as shown in Figure 5 (left), receiving a reward based on an estimate of the caption quality. After collecting a full batch of rewards, a RL optimization step is applied as shown in Figure 5 (right), and this process repeats iteratively until convergence.
We use the following notation: Let $T$ and $I$ be the sets of possible texts and images, with joint distribution $X$. Given image $i\in I$, an image captioning model $M$ with weights $\theta$ induces a conditional probability distribution $\pi_{\theta}(\cdot|\cdot)$ over generated captions $\hat{c}\in T$ conditioned on images $i\in I$. In the RL context, we refer to $\pi_{\theta}$ as the policy. A reward function $r:T\times T\times I\to\mathbb{R}$ assigns reward (or score) $r(\hat{c};c,i)$ to generated caption $\hat{c}$ relative to ground-truth caption $c$ and image $i$.
3.1 Reward Function
We wish to optimize for the competing objectives of output fidelity (low hallucination rate) and adequacy (including sufficient details to describe the input image), as optimizing for one of these alone causes the other to deteriorate (as shown in our ablations). We also wish to preserve other desired generation properties such as fluency and diversity. To achieve this, we design a reward function combining multiple objectives as follows:
Fidelity Objective. ($r_{f}$). To measure output fidelity to the input image, we use the GT reference captions as a proxy, checking for logical consistency via a pretrained Natural Language Inference (NLI) model. This outputs the probability $\overline{p}(\hat{c},c)$ that the generated text $\hat{c}$ logically contradicts $c$, serving as a strong signal for fidelity, as details which contradict ground-truth information about the image are guaranteed to be hallucinations. We scale to the range $[-1,1]$ by using $r_{f}(\hat{c};c):=1-2\overline{p}(\hat{c},c)$ as the fidelity reward. We implement this with BART Lewis et al. (2019) fine-tuned on the MNLI dataset Williams et al. (2018). We average values over all reference captions.
Adequacy Objective. ($r_{a}$). To measure adequacy (whether the output caption contains sufficient detail), we use BERTScore Zhang et al. (2019), a pretrained model measuring text quality relative to ground-truth references. We calculate its F1 value, scaled scale to be approximately in the range $[-1,1]$ as described in the appendix.
KL Regularization. Following prior work Jaques et al. (2017, 2019); Ziegler et al. (2020); Stiennon et al. (2020); Ouyang et al. (2022), we add a Kullback–Leibler (KL) divergence penalty to the reward model which constrains the agent to stay close to its initial policy $\pi_{0}$. This serves to prevent mode collapse (i.e. preserving diversity of outputs) and adversarial policies which over-optimize the reward function. The KL penalty adds a term proportional to $K(\hat{c};i):=-\log(\pi_{\theta}(\hat{c}|i)/\pi_{0}(\hat{c}|i))$ to the reward, which limits the agent from excessively distancing itself from the initial policy.
Combined Objective. Our total reward function takes the form $r(\hat{c};c,i):=\alpha\cdot r_{f}(\hat{c};c)+(1-\alpha)\cdot r_{a}(\hat{c};c)+\beta K(\hat{c};i)$, where $\alpha\in[0,1]$ and $\beta>0$ control the trade-off between objectives.
3.2 Learning Procedure
To optimize for caption generations that satisfy the desired properties (described above in Section 3.1), we adopt the Proximal Policy Optimization (PPO) RL algorithm Schulman et al. (2017), which has been used by recent works on text generation as discussed in Section 5. This is a policy gradient algorithm, meaning that it optimizes the parameters $\theta$ in order to (approximately) maximize the expected reward $L(\theta)=E_{i,c\sim X,\hat{c}\sim\pi_{\theta}(\hat{c}|i)}\left[r(\hat{c};c,i)\right]$. PPO extends the REINFORCE algorithm Sutton and Barto (2018), also known as SCST in the context of image captioning Rennie et al. (2017), by using a clipped surrogate objective to avoid instabilities.
4 Experiments and Results
OpenCHAIR Analysis. We analyze the utility of OpenCHAIR by comparing its distribution of objects to the existing closed-vocabulary CHAIR metric, as well as by performing a human evaluation to compare their correlations to human judgements of hallucinations.
[…]
We show that OpenCHAIR evaluations are grounded in human intuitions via a manual evaluation, comparing its performance to that of CHAIR. For each benchmark (OpenCHAIR and CHAIR), we generate captions for a random subset of its dataset and manually check object-level decisions (predicted as existing or hallucinated) for over 400 random objects. Results using various captioning models are found in Table 1. As the presence of hallucinations is highly imbalanced (the large majority of predicted objects are not hallucinated), we report balanced accuracy. We provide further details in the appendix, including full confusion matrices.
Surprisingly, although operating over a much more diverse scope, OpenCHAIR achieves higher accuracy than CHAIR. We identify that this stems from CHAIR’s heavy reliance on coarse synonym lists, as seen in Figure 6 (right). By assessing whether pairs of object names match using a knowledgeable LLM, OpenCHAIR performs finer-grained hallucination measurements and achieves superior accuracy even in the more general open-vocabulary setting. We note that this reflects a trade-off between true and false positives, as predicted objects may not be found in OpenCHAIR ground-truth lists despite being present in the accompanying images, due to the limited descriptive capacity of text used to generate images. See more details in the Appendix (Tables 3 and 4).
# Obj Types Balanced Accuracy
BLIP2 BLIP-L GIT-B OFA-L
CH 80* 0.844 0.774 0.899 0.810
OCH 2400 0.945 0.944 0.943 0.930
Table 1: Human Evaluation of OpenCHAIR and CHAIR. We perform a manual evaluation of OpenCHAIR and CHAIR object-level predictions, as described in Section 4. As seen above, OpenCHAIR covers a much larger variety of unique object types while also outperforming CHAIR in per-object predictive accuracy (of whether the given object is present or hallucinated). *CHAIR includes also a synonym list.
Figure 6: CHAIR Limitations. The left image exhibits CHAIR’s limited vocabulary. Out of all objects predicted by BLIP2, Scissors is the only object CHAIR considers during the evaluation. The right image illustrates a limitation stemming from CHAIR’s use of a fixed list of synonyms to coarsely aggregate different, semantically similar objects. Hallucinations that occur within the same synonym group are considered as a correct detection; in this example both Goose and Duck are defined as synonyms of Bird even though the image does not display a duck (but rather a goose).
Figure 7: Reducing Hallucinations While Maintaining Caption Quality. We show the relative improvement of state-of-the-art VLM models when optimized using MOCHa optimization on the COCO Caption Karpathy test set. CH and OCH refer to Chair and OpenCHAIR respectively. All results are generated by using their officially provided checkpoints and hyperparameters. Full numeric results are provided in the appendix.
Refer to caption Refer to caption Refer to caption
B A man in a suit and tie standing by another man in a suit and tie A person taking a tray of apples out of an oven A man sitting on a couch talking on a cell phone
B+M A man in a military uniform talking to a man in a suit and tie A person taking a pan of food out of an oven A man sitting on a couch using a laptop computer
Figure 8: Qualitative results of MOCHa applied to an image captioning model (BLIP-Large), along with baseline results without optimization (noted as B+M, B, respectively). We show captions (over COCO) produced from each model using beam search decoding with five beams. Hallucinated details are highlighted. The results illustrate that MOCHa encourages captions with high fidelity to the input image (avoiding hallucinations), while preserving a satisfying level of detail.
Figure 9: Fidelity-Adequacy graphs for pretrained (“initial”) and MOCHa-optimized BLIP models. As seen above, varying the reward weighting $\alpha$ adjusts the trade-off between caption fidelity (x-axis) and adequacy (y-axis), with intermediate values outperforming the initial model (“Initial”). This holds both for metrics we directly optimize (left) and additional metrics (right), illustrating the generalization ability of our approach.
As OpenCHAIR was produced by automatic generation followed by manual filtering, we investigate the effect of the small proportion of erroneous data removed (3%) on performance. Table 12 (appendix) shows that it only marginally impacts the resulting OpenCHAIR score, validating the high quality of its automatic generation mechanism.
MOCHa Implementation Details. We test image captioning with MOCHa on various SOTA image captioning models of varying architectures and across various sizes. In particular, we test BLIP Li et al. (2022a), BLIP-2 Li et al. (2023a) and GIT Wang et al. (2022). Following standard practice in RL-based image captioning, we use models that have first been fine-tuned on with a standard language modeling loss on the captioning dataset, and then applying PPO reinforcement with our reward function ($\alpha=0.5$). See the appendix for model checkpoints, parameter counts, and further training settings and hyperparameters.
We test our method on the MS-COCO Lin et al. (2015) captioning benchmark, using the data split of Karpathy and Fei-Fei Karpathy and Fei-Fei (2015) (113K items for training, 5K for evaluation). We report standard captioning metrics along with CHAIR Rohrbach et al. (2018) and OpenCHAIR over generated captions (beam search decoding with 5 beams). We also provide NLI ($\overline{p}$) and BERTScore values, directly optimized by MOCHa, as described in Section 3.1. In the appendix, we provide results on additional captioning datasets and metrics to further demonstrate generalization.
Quality Hallucination
Closed Open
[…]
BLIP 41.5 138.4 2.3 3.5 19.2 0.244
BLIP+L 5.5 0.0 12.1 35.4 31.8 0.321
BLIP+T 41.3 137.4 1.9 2.8 19.2 0.241
BLIP+M 41.9 139.6 2.1 3.1 18.3 0.206
BLIP-2 43.4 144.3 1.7 2.6 17.0 0.207
BLIP-2+L 5.7 0.0 12.1 33.6 28.4 0.259
BLIP-2+T 43.3 143.5 1.3 2.0 17.0 0.206
BLIP-2+M 44.0 144.3 1.4 2.3 16.6 0.199
Table 2: Comparison To Prior Works. Measured for BLIP-Large and BLIP-2. +L/T/M refer to LURE, TLC-A, and MOCHa respectively. B@4, C, CH, OCH, and $\overline{p}$ denote BLEU-4, CIDEr, CHAIR, OpenCHAIR, and NLI $p(\text{contr.})$. ) metrics respectively. All metrics are measured over MS-COCO test set, except for OCH which is measured over our OpenCHAIR benchmark.
MOCHa Results. Figure 7 presents quantitative results of image captioning models on MS-COCO showing the relative improvement of optimizing the baseline SOTA captioning models with MOCHa. As shown there, MOCHa improves measures of hallucinations in image captioning while preserving or even enhancing standard measures of caption quality. We note that this is despite the fact that the trade-off between these qualities may degrade one or the other when using a sub-optimal reward weighting (see ablations below). Figure 8 provides qualitative examples, illustrating that the MOCHa-optimized model generates captions consistent with the image while preserving a satisfying level of detail, consistent with our numeric results.
Our quantitative results show that MOCHa improves performance over base captioning models by most measures, across model architectures and sizes – not only among metrics that we directly optimize but also among non-optimized metrics, measuring general caption quality (e.g. CIDEr), closed-vocabulary hallucinations (CHAIR) and open-vocabulary hallucinations (OpenCHAIR). Along with our qualitative observations, this justifies our holistic approach to reducing hallucinations without restriction to a closed object list.
MOCHa Comparisons. In Table 2 we compare MOCHa to LURE Zhou et al. (2024) and TLC-A Petryk et al. (2023), current SOTA methods addressing VLM hallucinations, applied to the same pretrained BLIP and BLIP-2 models. LURE fails in the pure image captioning setting as its training procedure encourages long-form, highly detailed outputs. While these are in-distribution for instruction-tuned VLMs, they represent an increase in hallucinations relative to concise captions, as well as an extreme deviation from the reference texts; thus it degrades performance across metrics when applied to captioning models such as BLIP and BLIP-2. Regarding TLC-A, as it targets the objects in the closed-vocabulary object list of CHAIR, it shows an expected advantage in this metric, but does not improve the open-vocabulary hallucination rate (measured by OpenCHAIR) and even degrades other measures of caption quality, contrasting with the overall improvement shown by our method. More details and results are provided in Appendix B.3, B.4 and C.4.
A number of prior works have proposed dedicated methods for reduced-hallucination image captioning, often using data modification or building multi-component pipelines applied to older vision-language backbones. In Table 8 (appendix), we provide a comparison between these methods and SOTA foundation VLMs applied as-is, reprodducing results for the dedicated methods UD-L Biten et al. (2021), CIIC Liu et al. (2022), and COSNET Li et al. (2022b). We find SOTA VLMs outperform these methods across all metrics, motivating our focus on optimization applied on top of modern foundation models.
Ablations. We ablate the components of our reward function, finding that optimizing for fidelity alone degrades general caption quality, while optimizing for adequacy alone fails to improve hallucinations. This is seen in Figure 9 where extreme values of $\alpha$ (0 or 1) correspond to the edges of the curves. Adjusting the parameter $\alpha$ controlling the trade-off between objectives traces a Pareto frontier which outperforms the base model, showing that joint optimization of these objectives has a synergistic effect. The effects of each reward function component are also illustrated qualitatively in Figure 2 (appendix); removing $r_{f}$ from the reward function leads to increased hallucinations, and removing $r_{a}$ leads to captions that do not contain sufficient details. We provide full numeric results in the appendix, as well as ablating the effect of our chosen RL algorithm and of the KL-Penalty in our reward.
Figure 10: VLM Caption Hallucination Taxonomy. We illustrate metrics (left) and algorithms (right) for quantifying and mitigating hallucinations in image-conditioned text generation. We propose an explicit metric for measuring open-vocabulary hallucinations (OpenCHAIR) and an open-vocabulary hallucination mitigation algorithm (MOCHa). We mark each algorithm with the automatic hallucination rate metric with which it is evaluated (GreenOpenCHAIR, Red – CHAIR). Further details are provided in Section 5.